in a survey, 63% of americans said they own an anawering machine. If 12 Americans are selected at random, what is the probability that at least 10 own an answering machine.
Answer #1

X ~ Binomial (n,p)

Where n = 12 , p = 0.63

P(X) = nCx px ( 1 - p)n-x

So,

P( X >= 10) = P (X = 10) + P (X = 11) + P (X = 12)

= 12C10 0.6310 0.372 + 12C11 0.6311 0.371 + 12C12 0.6312 0.370

= 0.1205